# Local Conditional Controlling for Text-to-Image Diffusion Models

### v1

@@ Abstract
Diffusion models have exhibited impressive prowess in the text-to-image task. Recent methods add image-level controls, e.g., edge and depth maps, to manipulate the generation process together with text prompts to obtain desired images. This controlling process is globally operated on the entire image, which limits the flexibility of control regions. In this paper, we introduce a new simple yet practical task setting: local control. It focuses on controlling specific local areas according to user-defined image conditions, where the rest areas are only conditioned by the original text prompt. This manner allows the users to flexibly control the image generation in a fine-grained way. However, it is non-trivial to achieve this goal. The naive manner of directly adding local conditions may lead to the local control dominance problem. To mitigate this problem, we propose a training-free method that leverages the updates of noised latents and parameters in the cross-attention map during the denosing process to promote concept generation in non-control areas. Moreover, we use feature mask constraints to mitigate the degradation of synthesized image quality caused by information differences inside and outside the local control area. Extensive experiments demonstrate that our method can synthesize high-quality images to the prompt under local control conditions. Code is available at https://github.com/YibooZhao/Local-Control.
Figure 1: Previous image-level control and the proposed local control. Local control takes text prompt, image conditions and user-specified local control regions as input, generating images that are faithful to both the prompt and local control conditions.
@@ 1 Introduction
Large-scale text2image diffusion models [29, 27, 26] have showed remarkable capabilities in generating images conditioned on specified text prompts. However, text descriptions frequently fall short in efficiently or sufficiently conveying detailed control [38] over the final image generation results. Recent methods [44, 21] have introduced the incorporation of image-level spatial controls, such as edges, depth, and segmentation, into the text2image generation process. These methods operate across the entire image, ensuring that the final synthesized image retains the complete global structural information of the original. Such advancements have attracted considerable attention from both the academic and industrial sectors, as it facilitates a more controllable image synthesis process.
@@
Unfortunately, the conventional approach of enforcing control at the entire-image level lacks the fine-grained image manipulation capability that users often require. Considering the example in Figure 1, we might want to exclusively apply the cat canny condition. This naturally leads to a question: Can we confine our control to specific local areas using image conditions? In this paper, we specifically address this practical scenario and introduce the concept of local control as a novel and pragmatic paradigm. It focuses on controlling the generation of specific areas of an image according to user-defined local control conditions, while the rest of the image is conditioned by the original text prompt. As illustrated in Figure 1, only the canny condition of the cat is employed for controlling, while the remainder of the image is generated solely based on the provided text prompt. This novel manner empowers users with the ability to exert fine-grained and flexible control over the image generation process.
@@
An intuitive resolution for local control lies in directly adding local image conditions. However, this naive method leads to a catastrophic problem. The prompt concept that is most related to the local control condition dominates the generation process, while other prompt concepts are ignored. Consequently, the generated image cannot align with the input prompt. We call this problem local control dominance. We show the cases in the left column of Figure 2. Delving into the underlying mechanism, we find that the attention maps show high response areas for both the cat and dog within the local control regions, which impedes the generation of multiple objects, as seen here with the dog’s generation being hindered. It indicates that local control conditions force the model to focus on the controlled area and overlook other areas.
@@
To address this issue, we introduce an object regeneration strategy. It identifies the ignored objects in the cross attention map, enhancing them on the latent by adaptively operating attention scores under local and non-local regions. This manner guides the model to generate the most related objects to the image condition in the local region, meanwhile producing other desired objects in non-control regions for the text prompt. Additionally, we suppress attention scores for patches on the cross attention map which lacks the strongest response in the token dimension. It aims to facilitate the generation of objects in non-local region. Moreover, we use local region mask to constraint the feature output of Controlnet. This technology alleviate the degradation of synthesized image quality caused by differences in control information between non local and local control regions. All proposed strategies are operated at the inference stage. Overall, the proposed training-free approach achieves high-quality local control for both the text prompt and image conditions, without the need for any training or additional text-image paired datasets. Extensive experiments demonstrate the effectiveness of our method.
@@
  • We introduce a new task setting: local control, allowing for fine-grained controlling with image conditions. It aims to control specific regions within the image as desired by the user, while the remaining areas are conditioned solely by the original text prompt.
  • We pose the main challenge in local control, and propose a training-free method to resolve it. The proposed approach ensures the preservation of local control while maintaining the original generative capability of the diffusion models.
@@
Figure 2: Local control dominance. On the left, we exhibit a case of local control dominance, illustrated by the ControlNet cross-attention maps for “cat” and “dog”. The response for ”dog” in non-local regions is lower, leading to its absence in the generated image. On the right, we display the results of our method applied to local control with the same seed, where both “cat” and “dog” are successfully generated.
@@ 2.1 Text-to-Image Models
Early methods [41, 43] for text-to-image synthesis are limited in generating low-resolution images in specific domains. Extensive data volume[31] and abundant model capacity have led to significant progress in the field of text-to-image synthesis. In recent years, the diffusion model [3, 12] has achieved amazing results in text-to-image generation tasks and has gradually become a representative method in this field. Such as Imagen [29], Stable Diffusion [27], and DALL-E [26]. They leverage the denoising process, gradual reduction of noise to high-quality images through iterative optimization. In addition, the cross-attention mechanism incorporates content from text prompts into the generated images through CLIP[24] or large language models[25]. However, they are not sensitive to every concept in text prompts. To mitigate this problem, a few works like StructureDiffusion [4], Attend-and-Excite [2] and Multi-Concept T2I-Zero [36] refine the cross-attention map to strengthen all of the subjects to encourage the model to generate all concepts described in the text prompts.
2.2 Compositional Generation
Image composition involves placing a foreground object onto a background image. Simply placing a foreground object onto a background can lead to an unconvincing result. However, more recent works [42, 1, 20] utilize diffusion models to tackle the problems of fusion inconsistency, enhancing the efficiency of image composition. [1] takes an input image and a mask, and adjusts the area covered by the mask to match a guiding text prompt while leaving the unmasked regions unchanged. TF-ICON[20] is introduced as a pipeline that employs readily available diffusion models for cross-domain image-guided composition, eliminating the need for extra training, fine-tuning, or optimization. Paint by Example [42] trains a diffusion model to merge a reference image onto a source image at a location indicated by a mask of any shape, utilizing classifier-free guidance to assist the process. Recent advancements include VisorGPT[40], LayoutGPT [5], LLM blueprint[9], and ControlGPT[45], which are models specifically designed for planning visual layouts in image synthesis.
2.3 Controllable Generation
To facilitate controllable image generation, many works focus on customization concepts[33, 14, 8, 15], multi-concepts[17, 39, 30, 10], or conditional control[44, 21]. Text-driven control methods focus on adjusting prompts and manipulating token embedding. For instance, textual-inversion [7] injects the concept into the text embedding to control specific subject generation. MakeAScene [6] utilizes semantic masks to guide image generation. DreamBooth [28], Perfusion [34] and Custom Diffusion [16] allow for the personalization of content in the generated image through fine-tuning the image diffusion model using a small set of user-provided example images. Prompt2Prompt [11] and Plug-and-Play [35] explore the relationship between attention map and image representation. These methods can refine attention maps for image editing purposes. While they achieve high synthesis quality, text prompts may not always provide reliable structural guidance for the synthesis results. PITI [38] proposes to enhance structural guidance by reducing the gap between features derived from conditions. [37] suggests using the similarity gradient between the target sketch and intermediate model features to impose constraints on the structure of the final results. As concurrent works, [44] learns condition-specific ControlNet to enable guide generation for the pre-trained diffusion model. T2i-Adapter[21] introduces a simple and lightweight adapter to combine the internal knowledge in T2I models with external control signals. We find that existing controllable image generation methods typically focus on image-level control, which lack the ability to specify local control for generating images that align with both prompts and control conditions. Addressing this gap, we introduce local control as a novel and pragmatic paradigm. It enables precise control over specific image areas according to user-defined conditions, while the remaining parts are guided by the original text prompts. Additionally, we propose a seamless, training-free method compatible with existing ControlNet frameworks. Under the paradigm of local control, this method allows for high-quality image generation that closely aligns with the descriptions in the prompt.
Figure 3: Overview of our method. Given the text condition and local control condition, a latent variable $z_{T}$ is passed into the denoising network, i.e., UNet. At each denoising step, we apply Feature Mask Constraint (see Section 3.6) to the ControlNet branch output features. Cross-attention maps generated from U-net are refined by Focused Token Response (see Section 3.5), which suppresses weaker interfering responses to facilitate the concept generation of non-local regions. Furthermore, we use Object Regeneration (see Section 3.4) to update the resulting latent $z_{t}$ using attention score differences $l$, to identify and regenerate ignored objects in the cross-attention map.
@@ 3.1 Preliminaries
In this work, we apply our method on Stable Diffusion [27], which efficiently operates on the latent space of a pre-trained variational autoencoder. Specifically, given an input image $I$, the encoder $\mathcal{E}$ maps it to a latent $z$. image diffusion algorithms progressively add noise to the image and produce a noisy image $z_{t}$, where $t$ represents the number of times noise is added. Stable Diffusion enables the control of synthesized image content through additional textual inputs. The text prompt is first pre-processed to text tokens by the text encoder of pre-trained CLIP [22]. Diffusion model are trained using a squared error loss to learn a network $\epsilon_{\theta}$ that predicts the noise added to the noisy image $\bm{z}_{t}$:
$\mathcal{L}=\mathbb{E}_{\bm{z}_{0},\bm{t},\bm{c}_{t},,\epsilon\sim\mathcal{N}(0,1)}\Big{[}\|\epsilon-\epsilon_{\theta}(\bm{z}_{t},\bm{t},\bm{c}_{t}))\|_{2}^{2}\Big{]}$ (1)
@@
Relying solely on text prompts often proves insufficient for generating images that align with the user’s specific desires, recent methods integrate image-level spatial controls. ControlNet[44] utilizes the deep and robust encoding layers of a large-scale diffusion model. It employs zero convolutions that gradually increase the parameters from zero, ensuring that no detrimental noise impacts the fine-tuning process. During training, an additional spatial control condition $\bm{c}_{\text{f}}$ is required, and the training loss is:
$\mathcal{L}=\mathbb{E}_{\bm{z}_{0},\bm{t},\bm{c}_{t},\bm{c}_{\text{f}},\epsilon\sim\mathcal{N}(0,1)}\Big{[}\|\epsilon-\epsilon_{\theta}(\bm{z}_{t},\bm{t},\bm{c}_{t},\bm{c}_{\text{f}}))\|_{2}^{2}\Big{]}$ (2)
3.2 Overview
@@
In the context of local control, the spatial structure information required for the local control region is designated by the user as $\mathcal{C}_{f}$. Given a text prompt, the tokenizer produces a set of tokens $\mathcal{P}=\{\mathbf{p}_{i}\}$, and a set of corresponding spatial cross-attention maps $\mathcal{A}^{t}=\{\mathbf{A}^{t}_{i}\}$ in Unet can be accordingly obtained. The most straightforward method of enabling local control is to directly input $\mathcal{C}_{f}$ and $\mathcal{P}$ into the existing control methods. However, this approach leads to the issue of local control dominance (Figure 2).
In this paper, we propose a training-free method that synthesizes images using local control condition and the prompt as inputs. This method builds upon prior research [11] which has shown that the prominently responsive regions within the cross-attention map are crucial in determining the spatial positioning of generated objects. Our method is illustrated in Figure 3. Specifically, our method consists of three components: 1) Object Regeneration (see Section 3.4), which facilitates the generation of ignored objects; 2) Focused Token Response (see Section 3.5), which suppresses non-maximum attention scores for removing interference to facilitate the concept generation in non-local region; 3) Feature Mask Constraint (see Section 3.6), which mitigates image quality degradation caused by discrepancies in control information.
@@ 3.3 Control Concept Matching
where $\mathbf{M}$ denotes a binary spatial mask, which is generated from the local control region. It aims to selectively local control elements in the cross-attention maps. This approach ensures that the chosen object is optimally aligned with the user-defined control conditions.
Figure 4: Comparisons to ControlNet [44] with local image conditions.
3.4 Object Regeneration
Due to the impact of local control dominance, most objects often respond to the local control regions in cross-attention maps. An effective strategy to counter this effect involves directing the attention responses of objects in non-control regions away from the local control region. To achieve this, we introduce the object regeneration strategy, which is specifically designed to enhance the difference in response between the inside and outside of the local control region. We select a set of object tokens $\mathcal{S}=\{\mathbf{s}_{i}\}$ from the collection $\mathcal{P}$, where $\mathcal{P}$ represents the entire set of prompt tokens.
$\begin{split}L_{s_{i}}=D\big{[}&\max(G(\mathbf{A}_{s_{i}})\cdot\mathbf{M})\\ &-\max(G(\mathbf{A}_{s_{i}})\cdot(1-\mathbf{M}))\big{]}\end{split}$ (4)
When $s_{i}$ is the token corresponding to $\text{C}^{t}_{\text{control}}$, $D$ is set to -1; otherwise, it is 1. $G(\cdot)$ represents a Gaussian smoothing operator. $L_{s_{i}}$ uses the difference between the maximum attention scores inside and outside the local control region to represent distance.
$z_{t}=z_{t}-\alpha_{t}\nabla l$ (5)
$l$ is determined by $\max(L_{si})$ and is used to update the noised latent $z_{t}$ at the current time step $t$, where $\alpha_{t}$ is a scalar defining the step size of the gradient update. $l$ updates noised latents by increasing the attention scores that exhibit the maximum difference between patchs inside and outside the local control region. This approach directs the model to generate objects closely related to the control condition in the local region, while also creating objects in the non-control areas as given text prompt.

### v3

@@ Abstract
Diffusion models have exhibited impressive prowess in the text-to-image task. Recent methods add image-level structure controls, e.g., edge and depth maps, to manipulate the generation process together with text prompts to obtain desired images. This controlling process is globally operated on the entire image, which limits the flexibility of control regions. In this paper, we explore a novel and practical task setting: local control. It focuses on controlling specific local region according to user-defined image conditions, while the remaining regions are only conditioned by the original text prompt. However, it is non-trivial to achieve local conditional controlling. The naive manner of directly adding local conditions may lead to the local control dominance problem, which forces the model to focus on the controlled region and neglect object generation in other regions. To mitigate this problem, we propose Regional Discriminate Loss to update the noised latents, aiming at enhanced object generation in non-control regions. Furthermore, the proposed Focused Token Response suppresses weaker attention scores which lack the strongest response to enhance object distinction and reduce duplication. Lastly, we adopt Feature Mask Constraint to reduce quality degradation in images caused by information differences across the local control region. All proposed strategies are operated at the inference stage. Extensive experiments demonstrate that our method can synthesize high-quality images aligned with the text prompt under local control conditions.
Figure 1: Previous global control mechanism mainly synthesizes images similar to structure conditions, but has difficulty generating results aligned with text prompts. Even adding control mask could only produce concepts closest to the local condition. Therefore, we explore local control, which leverages text prompts, image conditions, and user-defined regions for local control as inputs. Our proposed method successfully generate images that are faithful to both the prompts and local control conditions.
@@ 1 Introduction
Large-scale text-to-image diffusion models (Saharia et al. 2022; Rombach et al. 2022; Ramesh et al. 2021) have showed remarkable capabilities in generating images conditioned on specified text prompts. However, text descriptions frequently fall short in sufficiently conveying detailed control (Wang et al. 2022) over the generation images. Recent methods (Zhang, Rao, and Agrawala 2023; Mou et al. 2023) have introduced the incorporation of image-level spatial controls, such as edges, depth, and segmentation, into the text-to-image generation process. These methods operate across the entire image, ensuring that the final synthesized image retains the complete global structural information of the original. Such advancements have attracted considerable attention from both the academic and industrial sectors, as it facilitates a more controllable image synthesis process.
Figure 2: In the ControlNet pipeline, local conditions dominate the image generation. The attention map for “dog” exhibits a higher response in local areas while a lower response in non-local regions where the dog should be generated. Meanwhile, we showcase the result of our method applied to local control , where both “cat” and “dog” are successfully generated.
@@
Unfortunately, the conventional approach of enforcing global control at the entire-image level lacks the fine-grained image manipulation capability that users often require. Considering the example in Figure 1, we might want to exclusively apply the cat canny condition, while allowing the generation in other regions fully leverages the capabilities of the text-to-image model. This naturally leads to a question: Can we confine our control to specific local region using image conditions? In this paper, we specifically address this practical scenario and introduce the concept of local control as a novel and pragmatic paradigm. It focuses on controlling the generation of specific regions of an image according to user-defined local control conditions, while the rest of the image is conditioned by the text prompt. As illustrated in Figure 1, only the canny condition of the cat is employed for controlling, while the remainder of the image is generated solely based on the provided text prompt. This novel manner empowers users with the ability to exert flexible control over the image generation process.
@@
An intuitive resolution for local control lies in directly adding local image conditions. However, this naive method leads to a catastrophic problem. The prompt concept that is most related to the local control condition dominates the generation process, while other prompt concepts are ignored. Consequently, the generated image cannot align with the input prompt. We call this problem local control dominance. We show the cases in the Figure 2. Delving into the underlying mechanism, we find that the attention map for the dog shows high response within the local control regions, which impedes the generation of multiple objects, as seen here with the dog’s generation being hindered. It indicates that local control conditions force the model to focus on the controlled region and overlook other regions.
@@
To address this issue, we introduce an object regeneration strategy. Specifically, we propose the Regional Discriminate loss to update the latents. It identifies the ignored objects in the cross attention map, enhancing them on the latent by adaptively operating attention scores under local and non-local regions. This manner guides the model to generate the most related objects to the image condition in the local region, meanwhile producing other desired objects in non-control regions for the text prompt. Additionally, we suppress attention scores for patches on the cross attention map which lack the strongest response in the token dimension. This technology alleviates token overlapping contribution to enhance object distinction and reduce duplication. Moreover, we use local region mask to constrain the feature outputs of ControlNet. This technology reduces the degradation of synthesized image quality caused by differences in control information between non-local and local control regions. All proposed strategies are operated at the inference stage. Overall, the proposed training-free approach achieves high-quality local control for both the text prompt and image conditions, without the need for any training or additional text-image paired datasets. Extensive experiments demonstrate the effectiveness of our method.
@@
  • We introduce a new task setting: local control, allowing for fine-grained controlling with image conditions. It aims to control specific regions within the image as desired by the user, while the remaining regions are conditioned solely by the original text prompt.
  • We pose the main challenge in local control, and propose a training-free method that includes three techniques to resolve it. The proposed method ensures the preservation of local control while maintaining the generative capability of the diffusion models.
@@ 2 Related Work
Text-to-Image Models. Early methods (Xu et al. 2018; Zhang et al. 2017) for text-to-image synthesis are limited in generating low-resolution images in specific domains.In recent years, extensive data volume LAION(Schuhmann et al. 2022) and abundant model capacity have led to significant progress in the field of text-to-image synthesis. The diffusion model (Dhariwal and Nichol 2021; Ho, Jain, and Abbeel 2020) has achieved amazing results in text-to-image generation tasks and has gradually become a representative method in this field. Such as Glide(Nichol et al. 2021), Imagen (Saharia et al. 2022), Stable Diffusion (Rombach et al. 2022), and DALL-E (Ramesh et al. 2021). They leverage the denoising process, gradual reduction of noise to high-quality images through iterative optimization. In addition, the cross-attention mechanism incorporates content from text prompts into the generated images through CLIP(Radford et al. 2021a) or large language models(Raffel et al. 2020).
Compositional Generation. Compositional Generation involves constructing or combining images semantically based on a set of given elements or attributes. (Yang et al. 2023; Avrahami, Lischinski, and Fried 2022; Lu, Liu, and Kong 2023) utilize diffusion models to tackle the problems of fusion inconsistency, enhancing the efficiency of image composition. Recent works focus on mutil-concept generation, Prompt2Prompt (Hertz et al. 2022) and Plug-and-Play (Tumanyan et al. 2023) explore the relationship between attention map and image representation. StructureDiffusion (Feng et al. 2022), Attend-and-Excite (Chefer et al. 2023) and Multi-Concept T2I-Zero (Tunanyan et al. 2023) refine the cross-attention map to strengthen all of the subjects to encourage the model to generate all concepts described in the text prompts. Gligen (Li et al. 2023b) achieves grounded text2img generation with text prompt and bounding box condition inputs. Boxdiff (Xie et al. 2023) propose a training-free method to control objects and contexts in the synthesized images. Unlike these methods, our approach aims to preserve the structural information of local regions during image generation, which facilitates finer-grained generation.
Controllable Generation. To facilitate controllable image generation, some works aim to provide reliable structural guidance for the synthesis results. PITI (Wang et al. 2022) proposes to enhance structural guidance by reducing the gap between features derived from conditions. (Voynov, Aberman, and Cohen-Or 2023) suggests using the similarity gradient between the target sketch and intermediate model features to impose constraints. ControlNet (Zhang, Rao, and Agrawala 2023) learns condition-specific network to enable guide generation for the pre-trained diffusion model. T2i-Adapter(Mou et al. 2023) introduces a lightweight adapter to combine the internal knowledge in T2I models with control signals. Uni-ControlNet(Zhao et al. 2024) proposes a condition injection strategy to learn an adapter for various controls, focusing on combined generation of multiple structure conditions. Recently, training-free approaches (Cao et al. 2023; Patashnik et al. 2023; Epstein et al. 2023; Mo et al. 2024) are proposed for controllable T2I diffusion. Although these existing structure-controllable image generation methods have achieved good control at the image-level, they lack the ability to provide localized control for generating images that align with both prompts and control conditions.
@@ 3 Method
In this section, we first introduce the preliminary knowledge required for local control. Secondly, we present Control Concept Matching which identifies objects for local control(see Section 3.2). We then detail our proposed training-free method that synthesizes images using the local control condition and the text prompt as inputs. Our method is illustrated in Figure 3. Specifically, our method consists of three components: 1) Regional Discriminate Loss (see Section 3.3), which updates the latents to facilitate the generation of ignored objects; 2) Focused Token Response (see Section 3.4), which suppresses non-maximum attention scores to enhance object distinction; 3) Feature Mask Constraint (see Section 3.5), which mitigates image quality degradation caused by discrepancies in control information.
@@ 3.1 Preliminaries
Stable Diffusion. We apply our method on Stable Diffusion (Rombach et al. 2022), which efficiently operates on the latent space. Specifically, given an input image $I$, the encoder maps it to a latent $z_{0}$. Diffusion algorithms progressively add noise $\epsilon$ to the $z_{0}$ and produce a noisy latent $z_{t}$, where $t$ represents the number of times noise is added. To achieve conditional generation, Stable Diffusion utilizes cross-attention between text embeddings $\mathbf{c}$ converted by CLIP (Radford et al. 2021b) and the UNet’s intermediate features $\varphi(\mathbf{x}_{t})$. The cross-attention map $\mathbf{A}$ can be defined as:
$\mathbf{A}=\text{Softmax}(\mathbf{Q}\mathbf{K}^{\top}/\sqrt{d}),$ (1)
where $\mathbf{Q}=\mathbf{W}_{Q}\varphi(\mathbf{x}_{t})$ and $\mathbf{K}=\mathbf{W}_{K}\mathbf{c}$ are the query and key matrices, respectively. Finally, $\epsilon_{\theta}$ can be trained by a MSE loss, which predicts the noise added to the latent $\bm{z}_{t}$:
$\mathcal{L}=\mathbb{E}_{\bm{z}_{0},\bm{t},\bm{c},\epsilon\sim\mathcal{N}(0,1)}\Big{[}\|\epsilon-\epsilon_{\theta}(\bm{z}_{t},\bm{t},\bm{c})\|_{2}^{2}\Big{]}$ (2)
@@
ControlNet. Relying solely on text prompts often fails to generate images that align with the user’s desires. ControlNet(Zhang, Rao, and Agrawala 2023) integrate global-level spatial controls which utilizes the encoding layers of the diffusion model. During training, an additional spatial control condition $\bm{c}_{\text{f}}$ is required, and the training loss is:
$\mathcal{L}_{c}=\mathbb{E}_{\bm{z}_{0},\bm{t},\bm{c},\bm{c}_{\text{f}},\epsilon\sim\mathcal{N}(0,1)}\Big{[}\|\epsilon-\epsilon_{\theta}(\bm{z}_{t},\bm{t},\bm{c},\bm{c}_{\text{f}}))\|_{2}^{2}\Big{]}.$ (3)
Therefore, ControlNet can utilize both text and image structure as conditions to control image generation.
@@
In the context of local control, given a text prompt, the tokenizer produces a set of tokens $\mathcal{P}=\{\mathbf{p}_{i}\}$, and a set of corresponding spatial cross-attention maps $\mathcal{A}^{t}=\{\mathbf{A}^{t}_{i}\}$ in UNet can be accordingly obtained. The most straightforward method of enabling local control is to directly input $\bm{c}_{\text{f}}$ and $\mathcal{P}$ into the existing control methods. However, this approach leads to the issue of local control dominance (Figure 2).
Figure 3: Overview of our method. Given the text condition and local control condition, a latent variable $z_{T}$ is passed into the denoising network, i.e., UNet. At each denoising step, we apply Feature Mask Constraint (see Section 3.5) to the ControlNet branch output features. Cross-attention maps generated from UNet are refined by Focused Token Response (see Section 3.4), which suppresses weaker interfering responses to enhance object distinction. Furthermore, we use Regional Discriminate Loss (see Section 3.3) to update the latent $z_{t}$, thereby identifying and regenerating ignored objects in the cross-attention map.
@@ 3.2 Control Concept Matching
where $\mathbf{M}$ denotes a binary spatial mask, which is generated from the local control region. It aims to selectively local control elements in the cross-attention maps. This approach ensures that the chosen object is optimally aligned with the user-defined local control conditions.
3.3 Regional Discriminate Loss
Due to the impact of local control dominance, most objects often respond to the local control regions in cross-attention maps. An effective strategy to counter this effect involves directing the attention responses of objects in non-control regions away from the local control region. To achieve this, we introduce the object regeneration strategy, which is specifically designed to enhance the difference in response between the inside and outside of the local control region.
Attend-and-excite(Chefer et al. 2023) indicates that the maximum value regions in the cross-attention map are key to defining the positioning of generated objects. Therefore, we choose the maximum value in the attention map to design a loss, named Regional Discriminate Loss (RDLoss). We adopt the max distance between the maximum values inside and outside the local region as the loss, so that each attention map can separate the objects in the local and non-local regions based on whether they are $\text{C}^{t}_{\text{control}}$ adaptive.
$\text{RDLoss}_{s_{i}}$ $=\mathbb{I}\left(\max(G(\mathbf{A}_{s_{i}})\cdot\mathbf{M})\right.$ (5)
$\quad\left.-\max(G(\mathbf{A}_{s_{i}})\cdot(1-\mathbf{M}))\right),$
we select a set of object token indices $\mathcal{S}=\{\mathbf{s}_{i}\}$ from the collection $\mathcal{P}$, where $\mathcal{P}$ represents the entire set of prompt tokens. $G(\cdot)$ represents a Gaussian smoothing operator. And the indicator function $\mathbb{I}$ is defined as follows:
$\mathbb{I}(s_{i}=\text{C}^{t}_{\text{control}})=\begin{cases}-1,&\text{if }s_{i}=\text{C}^{t}_{\text{control}},\\ 1,&\text{otherwise}.\end{cases}$ (6)
RDLoss uses the difference between the maximum attention scores inside and outside the local region to represent distance. In the attention maps for tokens not requiring local control conditions, we enhance the generation in non-local regions. Conversely, for tokens that require local control, we enhance generation in local regions. After defining the loss, we can update the latent in the following way:
$z_{t}=z_{t}-\alpha_{t}\nabla l,$ (7)
$l$ is determined by $\max(RDLoss_{si})$ in eq 5 and is used to update the noised latent $z_{t}$ at the current time step $t$, where $\alpha_{t}$ is a scalar defining the step size of the gradient update. $l$ updates noised latents by increasing the attention scores that exhibit the maximum difference between patchs inside and outside the local control region. This approach directs the model to generate objects closely related to the control condition in the local region, while also creating objects in the non-control regions as given text prompt.